# PEPT: Expert Finding Meets Personalized Pre-training

### v1

@@ Abstract.
Finding appropriate experts is essential in Community Question Answering (CQA) platforms as it enables the effective routing of questions to potential users who can provide relevant answers. The key is to personalized learning expert representations based on their historical answered questions, and accurately matching them with target questions. Recently, the application of Pre-trained Language Models (PLMs) have gained significant attraction due to their impressive capability to comprehend textual data, and are widespread used across various domains. There have been some preliminary works exploring the usability of PLMs in expert finding, such as pre-training expert or question representations. However, these models usually learn pure text representations of experts from histories, disregarding personalized and fine-grained expert modeling. For alleviating this, we present a personalized pre-training and fine-tuning paradigm, which could effectively learn expert interest and expertise simultaneously. Specifically, in our pre-training framework, we integrate historical answered questions of one expert with one target question, and regard it as a candidate aware expert-level input unit. Then, we fuse expert IDs into the pre-training for guiding the model to model personalized expert representations, which can help capture the unique characteristics and expertise of each individual expert. Additionally, in our pre-training task, we design: 1) a question-level masked language model task to learn the relatedness between histories, enabling the modeling of question-level expert interest; 2) a vote-oriented task to capture question-level expert expertise by predicting the vote score the expert would receive. Through our pre-training framework and tasks, our approach could holistically learn expert representations including interests and expertise. Our method has been extensively evaluated on six real-world CQA datasets, and the experimental results consistently demonstrate the superiority of our approach over competitive baseline methods.  [^1] This work is an extended and revised version of research that appeared at EMNLP 2022 (Peng and Liu, 2022).
Expert Finding, Personalization, Community Question Answering, Recommender Systems, Pre-trained Language Model
@@ 1. INTRODUCTION
Recently, the popularity of online CQA websites based on user-generated content has been continuously increasing (Zhu et al., 2011b; Srba and Bielikova, 2016). For instance, some large websites such as Zhihu, Stack Exchange, Quora provide convenient bridges to exchange knowledge and have successfully garnered a substantial user base (Hu et al., 2019; Yuan et al., 2020b; Hu et al., 2021; Li et al., 2022). These are websites where users can pose questions or offer responses related to topics of their interest or expertise (Fang et al., 2016; Peng et al., 2022a). Due to the large participation, some questions remain unanswered for extended periods, resulting in unsatisfactory response times for askers (Zhao et al., 2017; Yuan et al., 2020b). Therefore, routing questions to suitable experts who can provide satisfactory answers (Chang and Pal, 2013; Zhao et al., 2014) is crucial to CQA websites. Expert finding on the CQA websites can effectively recommend questions to suitable experts (Zhu et al., 2014) capable of providing high-quality answers. Then, the askers could receive answers quickly (Liu et al., 2022a). This approach has recently garnered considerable attention in both academic and industry.
@@
A key challenge in expert finding is the accurate matching of candidate experts with target questions (Li et al., 2019; Zhang et al., 2020). Thus, it is crucial to learn superiority expert representations to achieve precise question-expert matching. Existing expert finding methods mostly learn expert representations from limited supervised data (e.g., experts’ previously answered questions) and subsequently compute the expert-question matching scores. For example, Pal et al. (Pal et al., 2011) model users from their early historical activities, then use classification and ranking models to find potential experts. NeRank (Li et al., 2019) considers the question raiser’s profile and question content to learn entities representations with a network embedding method, which could help capture personalized expert representations. Peng et al. (Peng et al., 2022a) design a multi-view learning paradigm (PMEF) based on the question title, body and tag information. It could learn multi-view question/expert representations separately, and then compute question-expert relatedness for identifying experts.
@@
In recent years, the language modeling capability of PLMs (e.g., BERT (Devlin et al., 2018)), has demonstrated remarkable improvements. Pre-training the Transformer Encoder (Vaswani et al., 2017), on large unsupervised corpora enables models to learn more generalized knowledge representations. Extensive research has shown that pre-trained models are capable of capturing intricate sequence patterns from large amounts of raw data. Then it can transfer this knowledge to a range of downstream NLP tasks, especially in situations where limited training data is available (Brown et al., 2020; Raffel et al., 2020). Inspired by this, researchers been exploring the feasibility of pre-training technology for finding suitable experts (Liu et al., 2022b; Peng and Liu, 2022). For example, one preliminary work ExpertPLM (Peng and Liu, 2022) designs a tailored pre-training model for modeling expert based on the historical answered questions, which could help learn expert representations from unsupervised data.
@@
1) Interest Modeling. Experts’ interests reflect their willingness to answer questions, which could be captured from their historical answered questions. Existing Pre-trained Language Model (PLM)-based expert finding methods typically concatenate all questions an expert has answered and employ the Masked Language Modeling (MLM) task for modeling. However, this approach only captures expert interest based on word-level information and neglects question-level interest. In fact, expert interest is better represented by question-level semantics rather than individual keywords. Therefore, how to capture the question-level interest is important during pre-training.
@@
2) Expertise modeling. One way to assess an expert’s ability to answer questions is by using the vote scores, which indicate the level of CQA user appreciation for a particular answer. In other words, higher vote scores generally indicate higher levels of expertise associated with the user providing that answer. Thereby, to learn an expert’s fine-grained abilities of answering questions, we incorporate his/her historical answered question input embedding with corresponding encoded vote scores. Additionally, given one target question, we introduce a vote-oriented task to predict the vote score an expert would receive, empowering the model a better perception of an expert’s ability to answer the target question.
3) Personalized modeling. To further enhance personalized expert characteristics learning, we introduce the unique Expert ID into the model, which could indicate different expert personalized characteristics. One simplistic approach is to translate numerical expert IDs into their corresponding word embeddings, but this method can lead to excessive vocabulary size. Inspired by continuous/soft prompts (Lester et al., 2021), we directly input Expert ID vectors into the model by incorporating them with the expert-level information, which enhances the model to achieve personalized pre-training.
In a word, we propose to unify heterogeneous information into the pre-training framework to provide rich foundations for learning expert representation. Furthermore, we design a multi-task pre-training paradigm for guiding the model to capture fine-grained expert interest and expertise.
@@
In a word, our proposed model makes the following main contributions:
@@
  • We unify the previous questions, corresponding vote scores, and personalized information in the pre-training stage to model experts. Beneficial from this framework, we can directly conduct the expert finding based on the pre-trained model without gaps. Furthermore, during pre-training, we design a question-level MLM task and a vote-oriented task, which allows to better model the fine-grained expert characteristics.
@@ 2. RELATED WORK
Table 1. Characteristics of Different Expert Finding Methods.
@@ 2.1. Pre-training for Natural Language Processing
Recently, various of Pre-Training Models (PTMs) have demonstrated the efficacy of pre-training on large corpora and the capability to acquire general language knowledge, which can be effectively transferred to various downstream NLP tasks. (Qiu et al., 2020). Earlier word embedding methods, such as Word2Vec (Mikolov et al., 2013), Glove (Pennington et al., 2014) learned good word embeddings by capturing word co-occurrence information. These embeddings derived from these models could significantly improve various NLP missions (e.g., text classification (Lilleberg et al., 2015), named entity recognition (Lample et al., 2016)) by capturing the semantic meanings of words. However, these models are context-independent and predominantly trained using shallow models, which may restrict their effectiveness in specific scenarios that necessitate a more profound comprehension of language context.
To overcome these limitations, various of pre-training methods (Brown et al., 2020) based on the Transformer architecture (Vaswani et al., 2017) have been developed, showing their significant improvements in learning contextual language representations. These approaches, such as BERT (Devlin et al., 2018), MASS (Song et al., 2019), and BART (Lewis et al., 2020), mainly pre-train model parameters based on various neural network (e.g., transformer encoder, transformer decoder) and language modeling objectives. Then fine-tune the model parameters with downstream supervised data. By employing pre-training and fine-tuning paradigms, it can proficiently utilize extensive corpora to acquire comprehensive language knowledge. At the same time, the model could integrate task-specific information during the fine-tuning phase. Thus, these models have the potential to enhance the performance of diverse downstream tasks.
@@ 2.2. Pre-training for Recommender Systems
In recent times, pre-training technology has been increasingly used in recommendation models to enhance the performance of recommendation systems. Specifically, these methods (Kang and McAuley, 2018; Sun et al., 2019) leverage pre-training techniques to learn item co-occurrence information, which can be applied to improve the recommendation. For example, Cheng et al. (Cheng et al., 2021) proposed CLUE. It employed contrastive learning as its optimization objective during pre-training, with a focus on optimizing sequence-level representations. UniSRec (Hou et al., 2022) captured more transferable representations from the item texts for modeling users and items, which could effectively generalize to new recommendation scenarios. Then it fine-tuned the pre-trained model for conducting recommendation and user-oriented downstream tasks, which obtained better performance. Furthermore, many researchers have aimed to capture universal user embeddings that could be applied to various user-oriented downstream tasks (Yuan et al., 2020a; Xiao et al., 2021). For example, Xiao et al. (Xiao et al., 2021) proposed UPRec, which integrated heterogeneous user information and employed two pre-training tasks to learn user-aware representations.
@@ 2.3. Expert Finding
Traditional methods often relied on feature engineering (Robertson and Zaragoza, 2009; Cao et al., 2012; Pal et al., 2012) or topic modeling techniques (Riahi et al., 2012; Ji and Wang, 2013; Yang et al., 2013) to learn question and expert representations. Afterwards, the model utilized the question and expert representations for computing the relevance scores and routing questions to suitable experts. For instance, Zhou et al. (Zhou et al., 2012) captured local-/global-level characteristics based on questions, users, and relations between them, then these features were fed to the SVM (Smola and Schölkopf, 2004) for finding suitable experts. Chang et al. (Chang and Pal, 2013) employed LDA (Blei et al., 2003) to model the topic information from the question contents and took that information as the question representations for recommending experts. Neshati et al. (Neshati et al., 2017) introduced two learning algorithms that considered the temporal-aware aspects of expert finding. They combine factors such as topic similarity to predict a user’s likelihood of becoming an expert. Deep learning-based methods (Huang et al., 2013; Qiu and Huang, 2015; Wang and Nyberg, 2015) leveraged neural networks to model questions and experts, then computed the question-expert relevance for finding experts (Li et al., 2019; Fu et al., 2020; Ghasemi et al., 2021). For example, TCQR (Zhang et al., 2020) employed the pre-trained BERT to learn question semantic information and captured the answerer’s features based on the question semantic and temporal information for expert finding. Qian et al. (Qian et al., 2022) designed multi-perspective metapath-based network and considered vote scores for question routing, but they omitted the temporal interest of the expert. HQExpert employed (Liu et al., 2022c) different granularity semantic information and temporal interest to accurately identify experts and improve accuracy, but they could not consider personalized characteristics of the expert and the varying expert-question relationships.
@@ 3. PROBLEM DEFINITION
We first formal definition of the expert finding, then give some important notations and explanations used throughout the remainder of this paper.
In the expert finding task, the main goal is to match a target question, denoted as $q^{t}$, with the most appropriate expert who is capable of providing a relevant and accurate answer. Assuming that there is a candidate expert set $C^{u}=\{c^{u}_{1},\cdots,c^{u}_{M}\}$ and a target question $q^{t}$ to be answered respectively, where $M$ is the number of experts. In the set $C^{u}$, the candidate expert $c_{i}^{u}$ is associated with historically answered questions, denoted as $Q_{i}^{u}$. The $Q_{i}^{u}$ includes $\{q_{1},\cdots,q_{n}\}$, where $n$ is historical question number. Each question is represented as a question title including a sequence of words. The vote scores obtained for each previously answered question of the expert can be presented as $V^{u}_{i}=\{v_{1},\cdots,v_{n}\}$. It is important to note that, the individual who provides the “accepted answer” for a given question is considered the expert or the ground truth expert. Each question has only one “accepted answer”.
@@
$\mathcal{V}$ The vocabulary in pre-training
@@
$S_{v}$ Vote score predicted by the model
$\hat{S}_{v}$ Vote score ground truth

### v2

@@ Abstract.
Finding experts is essential in Community Question Answering (CQA) platforms as it enables the effective routing of questions to potential users who can provide relevant answers. The key is to personalized learning expert representations based on their historical answered questions, and accurately matching them with target questions. Recently, the application of Pre-trained Language Models (PLMs) have gained significant attraction due to their impressive capability to comprehend textual data, and are widespread used across various domains. There have been some preliminary works exploring the usability of PLMs in expert finding, such as pre-training expert or question representations. However, these models usually learn pure text representations of experts from histories, disregarding personalized and fine-grained expert modeling. For alleviating this, we present a personalized pre-training and fine-tuning paradigm, which could effectively learn expert interest and expertise simultaneously. Specifically, in our pre-training framework, we integrate historical answered questions of one expert with one target question, and regard it as a candidate aware expert-level input unit. Then, we fuse expert IDs into the pre-training for guiding the model to model personalized expert representations, which can help capture the unique characteristics and expertise of each individual expert. Additionally, in our pre-training task, we design: 1) a question-level masked language model task to learn the relatedness between histories, enabling the modeling of question-level expert interest; 2) a vote-oriented task to capture question-level expert expertise by predicting the vote score the expert would receive. Through our pre-training framework and tasks, our approach could holistically learn expert representations including interests and expertise. Our method has been extensively evaluated on six real-world CQA datasets, and the experimental results consistently demonstrate the superiority of our approach over competitive baseline methods.  [^1] This work is an extended and revised version of research that published in EMNLP 2022 (Peng and Liu, 2022).
Expert Finding, Personalization, Community Question Answering, Pre-trained Language Model
*Qiyao Peng and Hongyan Xu are equal contributions. Wenjun Wang is the corresponding author
@@ 1. INTRODUCTION
Recently, the popularity of online CQA websites based on user-generated content has been continuously increasing (Zhu et al., 2011b; Srba and Bielikova, 2016; Amendola et al., 2024a). For instance, some large websites such as Zhihu  [^2] https://www.zhihu.com/, StackExchange  [^3] https://stackexchange.com/, Quora  [^4] https://www.quora.com/ provide convenient bridges to exchange knowledge and have successfully garnered a substantial user base (Hu et al., 2019, 2021). These websites allow users to ask questions or provide responses related to topics of their interest or expertise (Fang et al., 2016; Peng et al., 2022a). Due to the large participation, some questions remain unanswered for extended periods, resulting in unsatisfactory response times for askers (Yuan et al., 2020b). Therefore, routing questions to suitable experts, who are capable of providing satisfactory answers, is crucial to CQA websites (Chang and Pal, 2013). Expert finding on the CQA websites can effectively recommend questions to suitable experts (Amendola et al., 2024b, c) capable of providing high-quality answers. Then, the askers could receive answers quickly (Liu et al., 2022a). This approach has recently garnered considerable attention in both academic and industry.
@@
A key challenge in expert finding is the accurate matching of candidate experts with target questions (Li et al., 2019; Zhang et al., 2020). Thus, it is crucial to learn personalized expert representations to achieve precise question-expert matching. Existing expert finding methods mostly learn expert representations from limited supervised data (e.g., experts’ previously answered questions) and subsequently compute the expert-question matching scores. For example, Pal et al. (Pal et al., 2011) model users from their early historical activities, then use classification and ranking models to find potential experts. NeRank (Li et al., 2019) considers the question raiser’s profile, question content and question answerer ID embedding to learn entity representations with a network embedding method, which could help capture personalized expert representations. Peng et al. (Peng et al., 2022a) design a multi-view learning paradigm to learn multi-view question representations based on the question title, body and tag information. Then the model introduces the expert ID information to personalized compute question-expert relatedness for identifying experts.
@@
In recent years, the language modeling capability of PLMs (e.g., BERT (Devlin et al., 2018)), has demonstrated remarkable improvements. Pre-training the Transformer Encoder model (Vaswani et al., 2017), on large unsupervised corpora enables models to learn more generalized knowledge representations. Extensive research has shown that pre-trained models are capable of capturing intricate sequence patterns from large amounts of raw data. These models can then transfer this knowledge to a range of downstream NLP tasks, especially in situations where limited training data is available (Brown et al., 2020; Raffel et al., 2020). Inspired by this, researchers have been exploring the feasibility of pre-training technology for finding suitable experts (Liu et al., 2022b; Peng and Liu, 2022). For example, one preliminary work, ExpertPLM (Peng and Liu, 2022) designs a tailored pre-training model for modeling expert based on the historical answered questions, which could help learn expert representations from unsupervised data.
@@
1) Interest Modeling. Experts’ interests can reflect their willingness to answer questions, which could be captured from their historical answered questions. Existing Pre-trained Language Model (PLM)-based expert finding methods typically concatenate all the questions an expert has answered. However, this approach only captures experts’ interests based on word-level information and neglects question-level interest. In fact, expert interest is better represented by question-level semantics rather than individual keywords. Therefore, how to capture the question-level interest is important during pre-training.
@@
2) Expertise modeling. One way to model an expert’s ability to answer questions is by using the vote scores, which indicate the level of CQA user appreciation for a particular answer. In other words, higher vote scores generally indicate higher levels of expertise associated with the user providing that answer. Thereby, to learn an expert’s fine-grained abilities of answering questions, we incorporate his/her historical answered question input embedding with corresponding encoded vote scores. Additionally, given one target question, we introduce a vote-oriented task to predict the vote score an expert would receive, empowering the model a better understanding of an expert’s ability to answer the target question.
3) Personalized modeling. To further enhance personalized expert characteristics learning, we introduce the unique Expert ID into the model, which could indicate different expert personalized characteristics. One simplistic approach is to translate numerical expert IDs into their corresponding word embeddings, but this method can lead to excessive vocabulary size. Inspired by continuous/soft prompts (Lester et al., 2021), we directly input Expert ID embeddings into the model by incorporating them with the expert-level information, which enhances the model to achieve personalized pre-training.
In a word, we propose to unify heterogeneous information into the pre-training framework to provide rich information for learning expert representation. Furthermore, we design a multi-task pre-training paradigm for guiding the model to capture fine-grained expert interest and expertise.
@@
Our proposed model has the following main contributions:
@@
  • We unify the historical questions, corresponding vote scores, and personalized information in the pre-training stage to model experts. Beneficial from the designing of this, the expert finding can be directly conducted based on the pre-trained model. Furthermore, during pre-training, we design a question-level MLM task and a vote-oriented task, which allows to better model fine-grained expert characteristics.
@@ 2. RELATED WORK
Table 1. Characteristics of Different Expert Finding Methods. The model possesses the feature marked with a check mark, while the feature marked with a cross is not possessed by the model.
@@ 2.1. Pre-training for Natural Language Processing
Recently, various of Pre-Training Models (PTMs) have demonstrated the efficacy of pre-training on large corpora and the capability to acquire general language knowledge, which can be effectively transferred to various downstream NLP tasks. (Yang et al., 2020, 2023; Bai et al., 2023). Earlier word embedding methods, such as Word2Vec , Glove learned good word embeddings by capturing word co-occurrence information. These embeddings derived from these models could significantly improve various NLP missions (e.g., text classification, named entity recognition) by capturing the semantic meanings of words. However, these models are context-independent and predominantly trained using shallow models, which may restrict their effectiveness in specific scenarios that necessitate a more profound comprehension of language context.
To overcome these limitations, various of pre-training methods (Brown et al., 2020; Wang et al., 2023; Sun et al., 2024) based on the Transformer architecture (Vaswani et al., 2017) have been developed, showing their significant improvements in learning contextual language representations. These approaches, such as BERT (Devlin et al., 2018), MASS (Song et al., 2019), and BART (Lewis et al., 2020), mainly pre-train model parameters based on various neural network (e.g., transformer encoder, transformer decoder) and language modeling objectives. Then fine-tune the model parameters with downstream supervised data. MASS (Song et al., 2019) is a transformer-based model that introduces the “masked target sequence reconstruction” task to address the limitation of traditional sequence-to-sequence models in machine translation. By employing pre-training and fine-tuning paradigms, it can proficiently utilize extensive corpora to acquire comprehensive language knowledge. At the same time, the model could integrate task-specific information during the fine-tuning phase. Thus, these models have the potential to enhance the performance of diverse downstream tasks.
@@ 2.2. Pre-training for Recommender Systems
In recent times, pre-training technology has been increasingly used in recommendation models to enhance the performance of recommendation systems. Specifically, these methods (Kang and McAuley, 2018; Sun et al., 2019) leverage pre-training techniques to learn item co-occurrence information, which can be applied to improve the recommendation. For example, SASRec (Kang and McAuley, 2018) employed a self-attention based sequential model to identify which items are “relevant” from a user’s action history for improving sequential recommendation. Sun et al. (Sun et al., 2019) learned item representations via fusing both the left and right context, then utilized the cloze objective to train the model. Cheng et al. (Cheng et al., 2021) proposed CLUE. It employed contrastive learning as its optimization objective during pre-training, with a focus on optimizing sequence-level representations. UniSRec (Hou et al., 2022) captured more transferable representations from the item texts for modeling users and items, which could effectively generalize to new recommendation scenarios. Then it fine-tuned the pre-trained model for conducting recommendation and user-oriented downstream tasks, which obtained better performance. Furthermore, many researchers have aimed to capture universal user embeddings that could be applied to various user-oriented downstream tasks (Yuan et al., 2020a; Xiao et al., 2021). For example, PeterRec (Yuan et al., 2020a) proposed a self-supervised model to learn user representations, which could be used to infer user profiles, such as the gender, age, preferences, etc. Xiao et al. (Xiao et al., 2021) proposed UPRec, which integrated heterogeneous user information and employed two pre-training tasks to learn user-aware representations.
@@ 2.3. Expert Finding
Traditional methods often relied on feature engineering (Robertson and Zaragoza, 2009; Cao et al., 2012; Pal et al., 2012) or topic modeling techniques (Riahi et al., 2012; Ji and Wang, 2013; Yang et al., 2013) to learn question and expert representations. Afterwards, the model utilized the question and expert representations for computing the relevance scores and routing questions to suitable experts. For instance, Zhou et al. (Zhou et al., 2012) captured local-/global-level characteristics based on questions, users, and relations between them, then these features were fed to the SVM (Smola and Schölkopf, 2004) for finding suitable experts. Chang et al. (Chang and Pal, 2013) employed LDA (Blei et al., 2003) to model the topic information from the question contents and took that information as the question representations for recommending experts. Neshati et al. (Neshati et al., 2017) introduced two learning algorithms that considered the temporal-aware aspects of expert finding. They combine factors such as topic similarity to predict a user’s likelihood of becoming an expert. Deep learning-based methods (Huang et al., 2013; Qiu and Huang, 2015; Wang and Nyberg, 2015) leveraged neural networks to model questions and experts, then computed the question-expert relevance for finding experts (Li et al., 2019; Fu et al., 2020; Ghasemi et al., 2021). For example, TCQR (Zhang et al., 2020) employed the pre-trained BERT to learn question semantic information and captured the answerer’s features based on the question semantic, temporal information and answerer embedding for expert finding. Qian et al. (Qian et al., 2022) designed a multi-perspective metapath-based network and considered vote scores for question routing, but they omitted the temporal interest of the expert. HQExpert employed (Liu et al., 2022c) different granularity semantic information and temporal interest to accurately identify experts and improve accuracy, but they could not consider personalized characteristics of the expert and the varying expert-question relationships. Vaibhav et al. (Krishna et al., 2022) proposed a personalized expert finding approach to consider the similarity between the content of the recent activity of a user (the “domain expertise”) and the content of a given new question. MPQR (Qian et al., 2022) utilized a heterogeneous information network (HIN) constructed from answering records and voting information to personalized learn multi-perspective representations of question answerers, question raisers, and questions. EFHM (Peng et al., 2022b) is a multi-grained (word-level, question-level and expert-level) hierarchical matching framework for expert finding in CQA websites, which could capture more fine-grained matching relationships and achieve superior performance. In the expert-level, the model introduce the expert ID embedding to personalized learn expert representations.
@@ 3. PROBLEM DEFINITION
We first formal definition of the expert finding, then give important notations and explanations used throughout the remainder of this paper.
In the expert finding task, the main goal is to match a target question, denoted as $q^{t}$, with the most appropriate expert who is capable of providing a relevant and accurate answer. Assuming that there is a candidate expert set $C^{u}=\{c^{u}_{1},\cdots,c^{u}_{M}\}$ and a target question $q^{t}$ to be answered respectively, where $M$ is the number of experts. In the set $C^{u}$, the candidate expert $c_{i}^{u}$ is associated with historically answered questions, denoted as $Q_{i}^{u}$. The $Q_{i}^{u}$ includes $\{q_{1},\cdots,q_{n}\}$, where $n$ is historical question number. Each question is represented as a question title including a sequence of words. The vote scores obtained for each previously answered question of the expert can be presented as $V^{u}_{i}=\{v_{1},\cdots,v_{n}\}$. It is important to note that, the answerer who provides the “accepted answer” for a given question is considered the expert, i.e., the ground truth. Each question has only one “accepted answer”.
@@
$\mathcal{V}_{w}$ The word vocabulary
$\mathcal{V}_{s}$ The vote score vocabulary